OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDAL Graph each pair of parametric equations on a scale from -10 to 10. x = 3 sin3t y = 3 cos3t

OpenStudy (anonymous):

Make t equal a value ftom -10 to 10 and plug into equation to find each x, y coordinate

OpenStudy (anonymous):

isnt there another way to do it? im not trying to be annoying i just dont know

OpenStudy (perl):

yes there is

OpenStudy (perl):

x^2 + y^2 = (3 cos t ) ^2 + (3sin t)^2 = 9cos^2 t + 9 sin^2 t = 9 ( cos^2 t + sin^2 t ) = 9 * 1 so we have a cartesian equation of a circle x^2 + y^2 = 9 it is centered at (0,0) and has radius 3